# Thermodynamic bounds on energy use in Deep Neural Networks

### v2

@@ Abstract
Note that in modern Mixture of Experts (MoE) architectures, only a small subset of network parameters is activated for each training sample. Thus, the effective number of active parameters $N_{a}$ may be much smaller than the total parameter count $N$ appearing in Eq. (16) for analog training. This suggests that, at least in principle, in silico training may outperform analog training, in sharp contrast to inference, where dynamic and quasi-static analog systems can be vastly more energy efficient than digital ones, as established by Eqs.(5) and (7).
Furthermore, digital systems offer an additional advantage: once trained, neural network models can be copied and deployed essentially for free—an operation that is highly nontrivial for physical systems trained via slow annealing. That being said, present-day digital computers still operate at least 7 orders of magnitude above Landauer’s limit, with no clear pathway for dramatically closing this gap. In contrast, the physical realization discussed in this work offers a plausible route to building systems capable of operating near the thermodynamic bounds on energy efficiency. In Table 1 we present a comparison between our results, Eqs. (5),(7) (16), and the actual energy use of the modern LLMs. We also include the respective estimates of minimal energy use by a digital computer, set by the Landauer limit.
@@
(model) Current     Landauer limit Analog bounds
@@
.2.2.2.2 $5\cdot 10^{-8}$ $5\cdot 10^{-9}$
Table 1: Comparison of our results with actual energy use by modern LLMs: LLama 65B (inference), and DeepSick V3 (training), as well as with the respective Landauer limits. The estimates are based on data from Refs. [32, 33], as well as official specifications of Nvidia GPUs A100 and H800. Analog bounds are given by Eqs. (5),(7) and (16).
Since the thermodynamic bounds determined in this work, Eq. (5) and (16) scale with $T$, one might anticipate that lowering the operating temperature would reduce the energy cost. While operating at a reduced temperature $T^{*}<T$ may offer practical benefits, it does not circumvent the fundamental limitations imposed by the Second Law of Thermodynamics. To demonstrate this, consider a physical DNN maintained at a temperature $T^{*}$ below the ambient temperature $T$. The energy $E$ used in operation must eventually be removed as heat. According to the Second Law, this removal requires performing work $W$ such that the overall entropy does not decrease:
@@
Finally, it is important to recognize that the conventional use of energy or work as thermodynamic ”currencies” is largely a legacy of the Industrial Revolution. Throughout this paper, where we referred to ”energy cost,” a more precise formulation would involve the appropriate version of free energy. In the context of the information age, however, a more natural and general measure of thermodynamic cost is negentropy—a concept promoted by Schrödinger in his book What is Life? [34] and later formalized by L. Brillouin [35, 36]. Negentropy can be defined as:
$J\equiv-\frac{\Delta S}{k_{B}\ln 2}=\frac{F}{k_{B}T\ln 2},$ (20)
where $F$ is the Helmholtz free energy of the system of interest, and $S$ is the total entropy, including the thermal bath. We have rescaled negentropy to ensure that it is measured in bits, same units Shannon’s information entropy $H$. Brillouin reinterpreted the Second Law of Thermodynamics as Information Principle, nearly a decade before Landauer’s work:
$\Delta(J+H)\leq 0,$ (21)
In the context of the present study, our key results, Eqs. (5), (7), and (16), should be reformulated as lower bounds on the negentropy required for inference and training in DNNs.
In conclusion, we have investigated fundamental thermodynamic bounds governing energy consumption in DNNs, during both inference and training phases. For dynamic analog systems, which can be broadly classified as neuromorphic, we concluded that the thermopdynamic bound on inference energy is primarily determined by the need to reset the neuron state after each cycle, leading to Eq. (5). We also demonstrated that quasi-static analog DNNs, described through an explicit Hamiltonian, can theoretically achieve reversible inference operations with no theoretical minimum of energy use imposed by thermodynamics. Training in quasi-static analog systems, however, is fundamentally constrained by a different thermodynamic bound, Eq. (16). This bound has the same scaling relationship with training dataset size $D$ and parameter number $N$ as the computing power needed for in silico training. The digital platforms retain certain pragmatic advantages, including the ability to optimize the training protocol by using MoE architecture, and the ease and negligible energy cost of duplicating trained models. Nevertheless, our analysis demonstrates that analog implementations hold significant promise for surpassing current digital systems in terms of practical energy efficiency.

### v3

@@ V Discussion
Note that in modern Mixture of Experts (MoE) architectures, only a small subset of network parameters is activated for each training sample. Thus, the effective number of active parameters $N_{a}$ may be much smaller than the total parameter count $N$ appearing in Eq. (17) for analog training. This suggests that, at least in principle, in silico training may outperform analog training, in sharp contrast to inference, where dynamic and quasi-static analog systems can be vastly more energy efficient than digital ones, as established by Eq. (5).
Furthermore, digital systems offer an additional advantage: once trained, neural network models can be copied and deployed essentially for free—an operation that is highly nontrivial for physical systems trained via slow annealing. That being said, present-day digital computers still operate at least 7 orders of magnitude above Landauer’s limit, with no clear pathway for dramatically closing this gap. In contrast, the physical realization discussed in this work may provide a plausible route to building systems capable of operating near the thermodynamic bounds. In Table 1 we present a comparison between our results, Eqs. (5),(17), and the actual energy use of the modern LLMs. We also include the respective estimates of minimal energy use by a digital computer, set by the Landauer limit.
@@
(model) Current     Landauer limit Analog bound
@@ 44 $5\cdot 10^{-8}$ 0
.2.2 $5\cdot 10^{-8}$ $5\cdot 10^{-9}$
Table 1: Comparison of our results with actual energy use by modern LLMs: LLama 65B (inference), and DeepSick V3 (training), as well as with the respective Landauer limits. The estimates are based on data from Refs. [26, 27], as well as official specifications of Nvidia GPUs A100 and H800. Analog bounds are given by Eqs. (5) and (17).
Figure 2: Illustration of the impossibility of beating the thermodynamic bound by cooling the physical DNN.
Since the thermodynamic bound on training energy (17) scales with $T$, one might anticipate that lowering the operating temperature would reduce the energy cost. While operating at a reduced temperature $T^{*}<T$ may offer practical benefits, it does not circumvent the fundamental limitations imposed by the Second Law of Thermodynamics. To demonstrate this, consider a physical DNN maintained at a temperature $T^{*}$ below the ambient temperature $T$. The energy $E$ used in operation must eventually be removed as heat. According to the Second Law, this removal requires performing work $W$ such that the overall entropy does not decrease:
@@
VI Conclusions
We have developed a thermodynamic framework for quasi-static analog DNNs by mapping them onto a physical system described by free energy $F$, Eq. (3). This approach provides a direct statistical mechanical interpretation of neural computation. Inference in such systems corresponds to quasi-static relaxation of neuron variables to a unique global minimum of the free energy. As a result, inference can proceed in a thermodynamically reversible manner, with no entropy production and zero minimal energy cost, Eq. 5). This radically distinguishes the quasi-static analog platform from digital hardware, where the nessesary multistability of bits leads to Landauer’s limit.
Importantly, for any fixed input the minimum satisfies $F_{\min}=0$ (up to a constant fluctuational correction, Eq.(4)), which means all constraints can be satisfied simultaneously without residual stress. This observation places quasi-static inference in close analogy with isostatic elastic networks, where the number of constraints matches the available degrees of freedom, allowing the system to accommodate them without storing elastic energy. By contrast, the training overconstrains the system, as both inputs and target outputs are clamped simultaneously. In this case, the network cannot reach $F=0$; instead, finite stresses $\sigma_{i}$ appear in the free-energy landscape [Eqs. (5)–(7)]. These stresses propagate backward through the architecture and serve as physical analogs of the error signals in backpropagation. When parameters such as weights and biases evolve slowly compared to neuron relaxation, their annealing naturally relaxes stresses, thereby reducing error. In this way, training arises intrinsically from the physics of the system, without the need for an explicit externally defined loss function. This feature has profound implications for the physical learning community: it demonstrates that learning in analog substrates can be realized through pure parameter annealing, guided only by thermodynamic relaxation.
At the same time, our analysis shows that such training is irreducibly dissipative. The lower bound on training energy, Eq. (17), scales extensively with both the number of parameters $N$ and the dataset size $D$. This identifies a fundamental asymmetry: inference can, in principle, be reversible, but adaptation through learning necessarily incurs finite thermodynamic cost. The structure of Eq. (17) also draws a striking parallel with thermodynamic uncertainty relations, extending their scope to learning in artificial neural systems.
In summary, the free-energy model provides a unifying statistical-mechanical foundation for DNNs, clarifying when energy-efficient analog computation is possible and when thermodynamics imposes unavoidable limits. It highlights quasi-static analog inference as a plausible route toward reversible computing, while also establishing that training, although elegantly realized through intrinsic parameter annealing, carries a fundamental energetic price. For the design of future physical learning machines, these insights suggest a path toward architectures where computation and adaptation emerge not from imposed algorithms, but from the thermodynamics of relaxation itself.